# Pair Counting without Binning - A New Approach to Correlation Functions in Clustering Statistics

### v1

@@ Abstract
This paper presents a novel perspective on correlation functions in the clustering analysis of the large-scale structure of the universe. We begin with the recognition that pair counting in bins of radial separation is equivalent to evaluating counts-in-cells (CIC), which can be modelled using a filtered density field with a binning-window function. This insight leads to an in situ expression for the two-point correlation function (2PCF). Namely, the spatial autocorrelation at different positions with a given separation can be equivalently interpreted as the cross-correlation between the original density field and its filtered one by the binning window function — at the same position. Essentially, the core idea underlying our method is to introduce a window function to define the binning scheme, enabling pair-counting without binning. This approach develops a concept of generalised 2PCF, which extends beyond conventional discrete pair counting by accommodating non-sharp-edged window functions. In the context of the multiresolution analysis for cosmic statistics (MRACS, Feng 2007), we can implement a fast algorithm for estimating generalised 2PCF. To extend this framework to N-point correlation functions (NPCF) using current optimal edge-corrected estimators, we developed a binning scheme that is independent of the specific parameterisation of polyhedral configurations. In particular, we demonstrate a fast algorithm for the three-point correlation function (3PCF), where triplet counting is accomplished by assigning either a spherical tophat or a Gaussian filter to each vertex of triangles. Additionally, we derive analytical expressions for the 3PCF using a multipole expansion in Legendre polynomials, accounting for filtered field (binning) corrections. Numerical tests using several suites of N-body simulation samples show that our approach aligns remarkably well with the theoretical predictions. Our method provides an exact solution for quantifying binning effects in practical measurements and offers a high-speed algorithm, enabling high-order clustering analysis in extremely large datasets from ongoing and upcoming surveys such as Euclid, LSST, and DESI.
@@ 1 Introduction
The large-scale inhomogeneities of cosmic density fields are seeded by primordial quantum fluctuations in the very early universe, which are predicted to be Gaussian distributed in the standard inflationary scenario (Guth & Pi, 1982; Hawking, 1982). Statistically, a Gaussian random field is fully characterised by its mean and second-order statistics, including the 2PCF and its Fourier counterpart, the power spectrum; these are the fundamental clustering indicators of the large-scale distribution of mass traced by galaxies (e.g. Peebles, 1980; Hamilton, 1988; Hawkins et al., 2003; Yang et al., 2003; Eisenstein et al., 2005a; Li et al., 2006; Zehavi et al., 2011; Wechsler & Tinker, 2018; DESI Collaboration et al., 2024, and references therein) and place strong constraints on the cosmological models (e.g. Cole et al., 2005; Eisenstein et al., 2005b; Tegmark et al., 2006; Blake et al., 2011; Beutler et al., 2011; Percival et al., 2010; Anderson et al., 2014; Hildebrandt et al., 2016; Alam et al., 2017; Ivanov et al., 2020; Alam et al., 2021a). However, driven by gravitational instability, the subsequent growth of density perturbations leads to highly nonlinear clustering on large scales, displaying significant non-Gaussian features. These non-Gaussianities cause information originally encoded in 2-point statistics to leak and cascade into a hierarchy of correlation functions at higher orders. The lowest order of non-Gaussian statistics is the 3PCF (the Fourier transform of which is the bispectrum), which provides complementary information about the background cosmology (e.g. Gagrani & Samushia, 2017; Agarwal et al., 2021; Alam et al., 2021a; Gualdi et al., 2021; Samushia et al., 2021; Novell-Masot et al., 2023) and an independent probe to models of structure formation (e.g. Jing & Börner, 1998; Gaztanaga & Frieman, 1994; Frieman & Gaztanaga, 1994; Scoccimarro et al., 2001; Jing & Börner, 2004; Kayo et al., 2004; Gaztañaga & Scoccimarro, 2005; Pan & Szapudi, 2005b; Nichol et al., 2006; Kulkarni et al., 2007a; Marín et al., 2008; Marín, 2011; McBride et al., 2011; Marín et al., 2013; Gil-Marín et al., 2015; Gil-Marín et al., 2017; Guo et al., 2015; Slepian et al., 2017; Pearson & Samushia, 2018; Slepian & Eisenstein, 2018; Veropalumbo et al., 2021; Sugiyama et al., 2023).
@@
The early motivation of this paper is to develop a fast algorithm for NPCF integrated into the Hermes toolkit (HypER-speed MultirEsolution cosmic Statistics) - an open source parallel/GPU-accelerated tool for cosmic statistics in Python (Feng et.al.,, 2024). The core algorithm implemented in Hermes is updated from the MRACS (Multi-Resolution Analysis for Cosmic Statistics) scheme(Feng, 2007), in which a spatial point process is approximated by a continuous distribution decomposed in terms of a set of compact basis functions, and the CIC distribution can be produced by convolution of the reconstructed density field with a low-pass window function specified by the given geometric volume.
@@
Since CIC, binning, and filtering share the same mathematical operation, a binning scheme can be converted to a filtered operation on the original density field. Unlike traditional binning schemes, the filters can be independent of the parameterisation of the polyhedral configuration. Consequently, NPCFs of arbitrary spatial configurations can be efficiently estimated by filtered fields, i.e., by binned density fields (Slepian & Eisenstein, 2018). In the sense of traditional counting, it is a counting without binning. The expense of dropping off binning in this way is that one needs to take the smoothing effects into account in theoretical modelling. We show that at least for 2/3PCFs of the filter field, the model calculation is not more intimidating than the one accounting for binning effects. Compared with binned correlation functions of the discrete sample, correlation functions of the filter field are particularly useful for investigating features at large scales. The cumulant correlators (Szapudi & Szalay, 1997; Szapudi, 1998; Munshi et al., 2000) are actually special cases of the correlation functions of the filtered field.
@@ 2.1 Quantifying the binning effect in 2PCF
[…]
@@
[…]
@@ 3.1 The MRACS algorithm for CIC
In the preceding subsection, we made an algebraic extension of pair-counting from arithmetic pair-counting with top-hat windows to general continuous functions and further introduced the concept of generalised 2PCF. Here, we describe briefly a numerical scheme for performing a fast computation based on the MRCAS algorithm (Feng, 2007) implemented in the Hermes toolkit (Feng et.al.,, 2024).
@@
where $w_{i}$ is an individual weight assigned to each particle, either a discrete or (truncated) continuous variable. For a galaxy sample, the weight could be a selection function (e.g. Singh 2021; Karim et al. 2023), or a sort of mark related to the intrinsic properties of galaxies and their environments et al. as introduced in various marked clustering statistics (e.g. Sheth & Tormen 2004; Sheth 2005; Skibba et al. 2006; White & Padmanabhan 2009; Simpson et al. 2011; Simpson et al. 2013; Skibba et al. 2013; White 2016; Pujol et al. 2017; Hernández-Aguayo et al. 2018; Armijo et al. 2018; Neyrinck et al. 2018; Valogiannis & Bean 2018; Satpathy et al. 2019; Philcox et al. 2020; Alam et al. 2021b; Massara et al. 2021; Xiao et al. 2022). In general, the number density could be generalised to any physical fields associated with the spatial process, e.g. velocity or momentum field.
@@ 3.2 Fast algorithm for 2PCF: an in-situ approach
where the orthogonality of basis functions has been used. According to Eq. (36), the pair-counting is simplified by a scalar product of the scaling function coefficients of two fields. Moreover, if we are only interested in measuring the 2PCF, the sum above can be performed in the wavenumber space, as ensured by Parseval’s theorem, which only requires performing the FFT once on the scaling coefficients $\epsilon_{j{\bf l}}$.
@@ 5.1 Quantifying the binning effect in 3PCF
$\hat{W}_{\text{sphere}}(|{\bf k}_{1}+{\bf k}_{2}|,R)=\sum_{m}(-)^{m}(2m+1)G_{m}(k_{1},k_{2})P_{m}(\hat{\bf k}_{1}\cdot\hat{\bf k}_{2}),$ (45)
@@
$P_{l}P_{m}=\sum_{n=|l-m|}^{l+m}\left(\begin{array}[]{lll}l&m&n\\ 0&0&0\end{array}\right)^{2}(2m+1)P_{m},$ (47)
@@
$C_{lmn}=(-1)^{m+n}(2m+1)(2n+1)\left(\begin{array}[]{lll}l&m&n\\ 0&0&0\end{array}\right)^{2}.$ (50)
@@ 6 Concluding Remarks
(1) According to the working definition of 2PCF based on various statistical estimators, we recognise that pair counting in bins is mathematically equivalent to evaluating CIC, which can be accomplished by convolving the density field with a window function. This insight yields an in-situ expression of 2PCF. Namely, the ex-situ 2-point autocorrelation function can be understood as an in situ cross-correlation between the original density field and its filtered one. The filter applied depends on the binning scheme in the pair-counting process. In this approach, the binning effect can be easily incorporated into theoretical models without leading to any ambiguities while comparing theoretical predictions with mock or observational data. Our approach extends the traditional concept of arithmetic pair counting, offering a generalised view of 2PCF with an arbitrarily chosen filter. It can be reasonably concluded that an optimal filter can be specifically designed so that the maximum information content from the survey data can be extracted.
@@ Acknowledgements
All simulation data analysed in this paper are available through the MultiDark database website https://www.cosmosim.org and the Quojote website https://quijote-simulations.readthedocs.io/en/latest/. The authors agree to make data supporting the results or analyses presented in their paper available upon reasonable request. The Hermes toolkit will be publicly available on GitHub together with the release of the relevant code paper Feng et.al., (2024) soon. An early test version of Hermes is also available from the corresponding author upon special request.
@@
$N_{\text{rot}}\sim 4\times 10^{4}\Bigl{(}\frac{n}{4.0\cdot 10^{-4}(h^{-1}\text{Mpc})^{-3}}\Bigr{)}^{2}\Bigl{(}\frac{R_{\text{max}}}{50h^{-1}\text{Mpc}}\Bigr{)}^{6}$ (60)

### v2

@@ Abstract
This paper presents a novel perspective on correlation functions in the clustering analysis of the large-scale structure of the universe. We begin with the recognition that pair counting in bins of radial separation is equivalent to evaluating counts-in-cells (CIC), which can be modelled using a filtered density field with a binning-window function. This insight leads to an in situ expression for the two-point correlation function (2PCF). Essentially, the core idea underlying our method is to introduce a window function to define the binning scheme, enabling pair-counting without binning. This approach develops an idea of generalised 2PCF, which extends beyond conventional discrete pair counting by accommodating non-sharp-edged window functions. In the context of multiresolution analysis, we can implement a fast algorithm to estimate the generalised 2PCF. To extend this framework to N-point correlation functions (NPCF) using current optimal edge-corrected estimators, we developed a binning scheme that is independent of the specific parameterisation of polyhedral configurations. In particular, we demonstrate a fast algorithm for the three-point correlation function (3PCF), where triplet counting is accomplished by assigning either a spherical tophat or a Gaussian filter to each vertex of triangles. Additionally, we derive analytical expressions for the 3PCF using a multipole expansion in Legendre polynomials, accounting for filtered field (binning) corrections. Our method provides an exact solution for quantifying binning effects in practical measurements and offers a high-speed algorithm, enabling high-order clustering analysis in extremely large datasets from ongoing and upcoming surveys such as Euclid, LSST, and DESI.
@@ 1 Introduction
The large-scale inhomogeneities of cosmic density fields are seeded by primordial quantum fluctuations in the very early universe, which are predicted to be Gaussian distributed in the standard inflationary scenario (Guth & Pi, 1982; Hawking, 1982). Statistically, a Gaussian random field is fully characterised by its mean and second-order statistics, including the 2PCF and its Fourier counterpart, the power spectrum; these are the fundamental clustering indicators of the large-scale distribution of mass traced by galaxies (e.g. Peebles, 1980; Hamilton, 1988; Hawkins et al., 2003; Yang et al., 2003; Eisenstein et al., 2005a; Li et al., 2006; Zehavi et al., 2011; Wechsler & Tinker, 2018; DESI Collaboration et al., 2024, and references therein) and place strong constraints on the cosmological models (e.g. Cole et al., 2005; Eisenstein et al., 2005b; Tegmark et al., 2006; Blake et al., 2011; Beutler et al., 2011; Percival et al., 2010; Anderson et al., 2014; Hildebrandt et al., 2016; Alam et al., 2017; Ivanov et al., 2020; Alam et al., 2021a). However, driven by gravitational instability, the subsequent growth of density perturbations leads to highly nonlinear clustering on large scales, displaying significant non-Gaussian features. These non-Gaussianities cause information originally encoded in 2-point statistics to leak and cascade into a hierarchy of correlation functions at higher orders. The lowest order of non-Gaussian statistics is the 3PCF (the Fourier transform of which is the bispectrum), which provides complementary information about the background cosmology (e.g. Gagrani & Samushia, 2017; Agarwal et al., 2021; Alam et al., 2021a; Gualdi et al., 2021; Samushia et al., 2021; Novell-Masot et al., 2023) and an independent probe to models of structure formation (e.g. Jing & Börner, 1998; Gaztanaga & Frieman, 1994; Frieman & Gaztanaga, 1994; Scoccimarro et al., 2001; Jing & Börner, 2004; Kayo et al., 2004; Gaztañaga & Scoccimarro, 2005; Pan & Szapudi, 2005b; Nichol et al., 2006; Kulkarni et al., 2007a; Marín et al., 2008; Marín, 2011; McBride et al., 2011; Marín et al., 2013; Gil-Marín et al., 2015, 2017; Guo et al., 2015; Slepian et al., 2017; Pearson & Samushia, 2018; Slepian & Eisenstein, 2018; Veropalumbo et al., 2021; Sugiyama et al., 2023).
@@
The early motivation of this paper is to develop a fast algorithm for NPCF integrated into the Hermes toolkit (HypER-speed MultirEsolution cosmic Statistics) - an open source parallel/GPU-accelerated tool for cosmic statistics in Python (Feng et al., in preparation). The core algorithm implemented in Hermes is updated from the MRACS (Multi-Resolution Analysis for Cosmic Statistics) scheme(Feng, 2007), in which a spatial point process is approximated by a continuous distribution decomposed in terms of a set of compact basis functions, and the CIC distribution can be produced by convolution of the reconstructed density field with a low-pass window function specified by the given geometric volume.
@@
Since CIC, binning, and filtering share the same mathematical operation, a binning scheme can be converted to a filtered operation on the original density field. Unlike traditional binning schemes, the filters can be independent of the parameterisation of the polyhedral configuration. Consequently, NPCFs of arbitrary spatial configurations can be efficiently estimated by filtered fields, i.e., by binned density fields (Slepian & Eisenstein, 2018). In the sense of traditional counting, it is a counting without binning. The expense of dropping off binning in this way is that one needs to take the smoothing effects into account in theoretical modelling. We show that at least for 2/3PCFs of the filtered field, the model calculation is not more intimidating than the one accounting for binning effects. Compared with binned correlation functions of the discrete sample, correlation functions of the filtered field are particularly useful for investigating features at large scales. The cumulant correlators (Szapudi & Szalay, 1997; Szapudi, 1998; Munshi et al., 2000) are actually special cases of the correlation functions of the filtered field.
@@ 2.1 Quantifying the binning effect in 2PCF
$n_{R,\Delta R}({\bf x}_{i})=\langle n({\bf x}_{i}+{\bf R})\rangle_{V_{R,\Delta R}}=\frac{1}{V_{R,\Delta R}}\int_{V_{R,\Delta R}}n({\bf x}_{i}+{\bf R})d^{3}{\bf R},$ (4)
@@
$n_{R,\Delta R}({\bf x})=W_{R,\Delta R}(\mathbf{x})\circ n({\bf x})=\int W_{R,\Delta R}(\mathbf{x-x}^{\prime})n({\bf x}^{\prime})d^{3}{\bf x}^{\prime}.$ (5)
@@ 2.2 Generalised 2PCF: from ex-situ to in-situ
It is noted that we have only considered the 2PCF in real space, assuming spherical symmetry, which also applies to the introduced binning window function. However, the generalised 2PCF allows for window functions of arbitrary forms, tailored to match the observed symmetry of the matter distribution. For example, the spherical symmetry is broken in the redshift space and the 2PCF $\xi(\pi,\sigma)$ is typically parameterised by both the radial (redshift) distance $\pi$ and the transverse distances $\sigma$, which correspond to a ring-structured binning window function. Within the framework of the generalised 2PCF, the binning window function with an axially symmetric shape can be optimally chosen to improve redshift distortion measurements (Li et al. in preparation).
@@ 3.1 The MRACS algorithm for CIC
In the preceding subsection, we made an algebraic extension of pair-counting from arithmetic pair-counting with top-hat windows to general continuous functions and further introduced the concept of generalised 2PCF. Here, we describe briefly a numerical scheme for performing a fast computation based on the MRCAS algorithm (Feng, 2007) implemented in the Hermes toolkit (Feng et al., in preparation).
@@
where $w_{i}$ is an individual weight assigned to each particle, either a discrete or (truncated) continuous variable. For a galaxy sample, the weight could be a selection function (e.g. Singh 2021; Karim et al. 2023), or a sort of mark related to the intrinsic properties of galaxies and their environments et al. as introduced in various marked clustering statistics (e.g. Sheth & Tormen 2004; Sheth 2005; Skibba et al. 2006; White & Padmanabhan 2009; Simpson et al. 2011, 2013; Skibba et al. 2013; White 2016; Pujol et al. 2017; Hernández-Aguayo et al. 2018; Armijo et al. 2018; Neyrinck et al. 2018; Valogiannis & Bean 2018; Satpathy et al. 2019; Philcox et al. 2020; Alam et al. 2021b; Massara et al. 2021; Xiao et al. 2022). In general, the number density could be generalised to any physical fields associated with the spatial process, e.g. velocity or momentum field.
@@ 3.2 Fast algorithm for 2PCF: an in-situ approach
where the orthogonality of basis functions has been used. According to Eq. (36), the pair-counting is simplified by a scalar product of the scaling function coefficients of two fields. Moreover, if we are only interested in measuring the 2PCF, the above sum can be performed in the wavenumber space, as ensured by Parseval’s theorem, which only requires performing the FFT once on the scaling coefficients $\epsilon_{j{\bf l}}$, regardless of the number of radial bins we applied. On other hand, since FFT calculations are computationally inexpensive on most standard platforms, if we wish to store the filtered density fields for other purposes, we will need to perform an additional inverse FFT calculation for each radial bin of interest.
@@ 5.1 Quantifying the binning effect in 3PCF
$\hat{W}_{\text{sphere}}(|{\bf k}_{1}+{\bf k}_{2}|,R)=\sum_{m}(-1)^{m}(2m+1)G_{m}(k_{1},k_{2},R)P_{m}(\hat{\bf k}_{1}\cdot\hat{\bf k}_{2})$ (45)
@@
$P_{l}P_{m}=\sum_{n=|l-m|}^{l+m}\left(\begin{array}[]{lll}l&m&n\\ 0&0&0\end{array}\right)^{2}{(2n+1)P_{n}},$ (47)
@@
$C_{lmn}=(-1)^{m+n}\frac{1}{4\pi}(2l+1)(2m+1)(2n+1)\left(\begin{array}[]{lll}l&m&n\\ 0&0&0\end{array}\right)^{2}.$ (50)
@@ 6 Concluding Remarks
(1) According to the working definition of 2PCF based on various statistical estimators, we recognise that pair counting in bins is mathematically equivalent to evaluating CIC, which can be accomplished by convolving the density field with a window function. This insight yields an in-situ expression of 2PCF. Namely, the ex-situ 2-point autocorrelation function can be understood as an in-situ cross-correlation between the original density field and its filtered one. The filter applied depends on the binning scheme in the pair-counting process. In this approach, the binning effect can be easily incorporated into theoretical models without leading to any ambiguities while comparing theoretical predictions with mock or observational data. It is emphasised here that sharp edged filters are necessary for pair counting in a discrete point sample; however, our approach is based on the reconstructed continuous density field and extends the traditional concept of arithmetic pair counting, offering a generalised view of 2PCF with an arbitrarily chosen filter. Of importance, various cosmological applications can benefit from the flexibility in the filter selection, especially with the availability of a rich variety of non-sharp edged filters to adapt to specific purpose. For example, in the forthcoming paper (Li et al., in preparation), we introduced two axial symmetric window functions - disk and cylinder shell - to extract cosmological information in redshift space, and the results show that it can significantly improve the constraints on cosmological parameter estimation. Thus, it can be reasonably concluded that an optimal filter can be specifically designed so that the maximum information content encoded in the survey data can be extracted, although much work remains to be done in this respect.
@@ Acknowledgements
All simulation data analysed in this paper are available through the MultiDark database website https://www.cosmosim.org and the Quojote website https://quijote-simulations.readthedocs.io/en/latest/. The authors agree to make data supporting the results or analyses presented in their paper available upon reasonable request. The Hermes toolkit will be publicly available on GitHub together with the release of the relevant code paper soon (Feng et al., in preparation). An early test version of Hermes is also available from the corresponding author upon special request.
@@
$N_{\text{rot}}\sim 4\times 10^{4}\Bigl{(}\frac{n}{4.0\times 10^{-4}(h^{-1}\text{Mpc})^{-3}}\Bigr{)}^{2}\Bigl{(}\frac{R_{\text{max}}}{50h^{-1}\text{Mpc}}\Bigr{)}^{6}$ (60)